I have the hardest time with word problems. For this problem I need to use the discriminate to determine the number of solutions of the quadratic equation, and whether the solutions are real or complex. It is not necessary to find the roots; just determine the number and types of solutions.

The hudson river flows at a rate of 3 miles per hour. A patrol boat travels 60 miles upriver, and returns in a total time of 9 hours. What is the speed of the boat in still water?

Answer by scott8148(6628)

time = distance / rate ___ let x ="speed in still water"

[60 / (x - 3)] + [60 / (x + 3)] = 9

multiplying by (x - 3)(x + 3) ___ 60x - 180 + 60 x + 180 = 9x^2 - 81

0 = 9x^2 - 120x - 81

0 = 3x^2 - 40x - 27

discriminant = (-40)^2 - (4 * 3 * (-27)) = 160 + 324